An auto battery company claims that their batteries' mean life is 50 months. In order to check this cla, a DTI researcher took a random sample 18 of these batteries and found that the mean life is 48.8 months with standard deviation of 7 months. Aasume the battery life follows a normal destribution, test with 90% confidence wether the companies'claim is different from the true mean.




1
Expert's answer
2022-06-06T05:21:52-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=50"

"H_1:\mu\not=50"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=17" and the critical value for a two-tailed test is "t_c =1.7396."

The rejection region for this two-tailed test is "R = \{t:|t|>1.7396\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{48.8-50}{7/\sqrt{18}}=-0.7273"


Since it is observed that "|t|=0.7273<1.7396=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=17" degrees of freedom, "t=-0.7273" is "p=0.47694," and since "p= 0.47694>0.10=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 50, at the "\alpha = 0.10" significance level.